Evaluate 1^2 + 1^3 + 2^2 + 2^3 + 3^2 + 3^3 + ... + 150^2 + 150^3
A. 128,384,358
B. 129,391,900
C. 135,391,800
D. 147,920,368

Answer by math_tutor2020(3817)


Hints:

1^2+2^2+3^2+...+n^2 = n(n+1)(2n+1)/6
1^3+2^3+3^3+...+n^3 = (n(n+1)/2)^2